### v1

# AI-driven View Guidance System in Intra-cardiac Echocardiography Imaging

Abstract
Intra-cardiac Echocardiography (ICE) is a crucial imaging modality used in electrophysiology (EP) and structural heart disease (SHD) interventions, providing real-time, high-resolution views from within the heart. Despite its advantages, effective manipulation of the ICE catheter requires significant expertise, which can lead to inconsistent outcomes, particularly among less experienced operators. To address this challenge, we propose an AI-driven closed-loop view guidance system with human-in-the-loop feedback, designed to assist users in navigating ICE imaging without requiring specialized knowledge. Our method models the relative position and orientation vectors between arbitrary views and clinically defined ICE views in a spatial coordinate system, guiding users on how to manipulate the ICE catheter to transition from the current view to the desired view over time. Operating in a closed-loop configuration, the system continuously predicts and updates the necessary catheter manipulations, ensuring seamless integration into existing clinical workflows. The effectiveness of the proposed system is demonstrated through a simulation-based evaluation, achieving an 89% success rate with the 6532 test dataset, highlighting its potential to improve the accuracy and efficiency of ICE imaging procedures.
{IEEEkeywords}
Intra-Cardiac Echocardiography (ICE), Ultrasound navigation, View guidance, Deep learning, Mamba
1 Introduction
[…]
The Intra-cardiac Echocardiography (ICE) is a sophisticated imaging modality that offers real-time, high-resolution views from within the heart, making it an invaluable tool in both electrophysiology (EP) and structural heart disease (SHD) interventions. The ICE is commonly used in procedures such as atrial fibrillation ablation, transcatheter valve repairs, and closures of septal defects. The key advantages of ICE include its ability to provide clear and detailed near-field images of cardiac structures, perform under conscious sedation, and be easily manipulated and interfaced with other interventional equipment. Additionally, the ICE has been shown to reduce procedural and fluoroscopy times, decrease overall radiation exposure to both the patient and physician, and shorten hospital stays. As a result, the ICE has become widely used in interventional procedures for the management of complex cardiac conditions [1, 2, 3, 4].
The primary use cases of the ICE imaging involve visualizing target anatomy, detecting and tracking therapeutic devices, and validating treatments in real-time. For ICE imaging to be effective, the ICE catheter must be accurately positioned at the target location. Achieving this requires physicians to have significant expertise in interpreting anatomical views via ICE images and skillfully maneuvering the ICE catheter using two knobs (anterior-posterior, right-left) and the rotating/translating the catheter body itself to the target anatomical views, all while primarily focusing on the therapeutic device manipulation [5, 6]. This complexity can be particularly challenging for those who are less experienced. As a result, the procedure is largely dependent on the skill of the operator, which might lead to inconsistent outcomes and extended procedure times, emphasizing the importance of experience in achieving optimal results.
Figure 1: Overview of the proposed view guidance system. When users want to navigate to clinically pre-defined views during the procedures, our proposed system provides continuous guidance on how to manipulate the ICE catheter via an interactive device (e.g., a touchpad for view selection and a feedback monitor) until the target view is achieved. (ACUSON Origin Cardiovascular Ultrasound System and AcuNav Lumos ICE catheter, Source: Siemens Healthineers)
In this research, we propose an AI-driven closed-loop view guidance system for ICE imaging, designed to enable even novice users to effectively utilize ICE imaging without specialized knowledge. Figure 1 presents an overview of the proposed view guidance system. Specifically, our proposed method models the relative position and orientation vectors between arbitrary views and clinically defined common ICE views [7, 8, 9] — such as Right Ventricle (RV), Left Ventricle (LV), Left Atrial Appendage (LAA), Left Pulmonary Vein (LPV), Right Pulmonary Vein (RPV), or Esophagus (ESO) views — within a unified spatial coordinate system. The users are guided on how to manipulate the ICE catheter to transition from the current view to the desired view over time. The proposed model operates in a closed-loop configuration with a human in the loop: it predicts the necessary manipulations to the ICE catheter’s state between the current and goal views, and estimates the current catheter state to manipulate by comparing the differences between the current and home views, as well as the goal and home views, as the catheter is manipulated. This proposed system avoids disrupting the existing clinical workflow, making it seamlessly integrable into current practices. Finally, the system is evaluated by simulation-based demonstrator, where the boundaries of the predictions from our quantile loss were assessed with the collected clinical data.
2 Related works and Background
2.1 US view navigation
To improve the clinical workflow, various research efforts have focused on developing guidance systems for tasks such as view finding and clinical decision-making. In particular, Artificial Intelligence (AI)-based methods have been widely utilized to guide and support these workflow. For instance, Narang et al. [10] proposed novel software that provides real-time prescriptive guidance to novice operators for obtaining the Trans-Thoracic Echocardiographic (TTE) images. Similarly, Sabo et al. [11] proposed a real-time guidance system to improve the echo-cardiographic acquisition process, and Pasdeloup et al. [12] developed a method to help users navigate toward preferred standard views during scanning. Additionally, Amadou et al. [13] demonstrated a synthetic image generation pipeline for navigation in an echocardiography view classification experiment. Building on this, Amadou et al. [14] proposed a framework using reinforcement learning to guide novice sonographers in navigating to standard interventional views during Transesophageal echocardiography (TEE).
While many pioneering works have significantly improved clinical workflows, the aforementioned studies primarily deal with acquiring echocardiograms externally or rely on simulation data. In contrast, our research is the first to specifically address a view guidance system for ICE catheter manipulation based on real clinical data. This approach requires a real-time understanding of internal cardiac structures, providing continuous guidance through the interactive mapping of images to spatial and catheter joint space. Furthermore, unlike simulation-based methods, which may not fully capture the complexities of real clinical environments, our method is designed for seamless integration into actual clinical workflows, effectively guiding the catheter within the heart’s intricate anatomy.
Although some robotic systems [15, 16, 17] exist to assist with catheter control, they are often expensive and complex to integrate into the clinical workflow, typically focusing on reproducing previously saved views by manual view survey [15, 16] or relying on positional sensors within a spatial coordinate system [17]. In contrast, our approach relies solely on ICE-imaging-based state estimation to guide the user in manipulating the ICE catheter within the control space from the current view to the target view, eliminating the need for positional sensors and avoiding disruption to the existing clinical workflow.
2.2 From Transformer to the Mamba and its medical applications
To address the relationship between non-linear behaviors in images and associated multi modal states, many recent studies have turned to AI, particularly leveraging a Transformer architecture [18]. The Transformer has been widely adopted in vision tasks due to its powerful ability to understand context across entire images through its self-attention mechanism, as well as its scalability to large models and robustness. As a result, many tasks that traditionally relied on CNN structures are now being addressed with Transformer-based architectures [19, 20, 21, 22, 23].
Recently, a new alternative to the transformer architecture, called Mamba [24], has emerged. Unlike the transformer architecture, which relies on self-attention mechanisms with quadratic computational complexity for long sequences, the State Space Model (SSM) maintains linear computational complexity. Mamba is based on the SSM, which can be represented as follows:
$\begin{split}h[k]&=\hat{A}h_{[}k-1]+\hat{B}x[k]\\ y[k]&=Ch[k],\end{split}$ (1)
where the $h[k]$ represents the state at the $k$-th step, and the matrix $C$ is the projection matrix that projects the current state to the output. The matrices $\hat{A}$ and $\hat{B}$ are the discretized versions of the system matrix and projection matrix, respectively, and they can be represented as follows:
$\begin{split}\hat{A}&=\exp(\Delta A)\\ \hat{B}&=(\Delta A)^{-1}(\exp(\Delta A)-I)\cdot\Delta B,\end{split}$ (2)
where the $\Delta$ denotes the timescale parameter. Mamba implements the Selective Scan Mechanism with SSM (S6), filtering out irrelevant information to effectively handle long-range sequences and induction heads. To achieve this, it allows the matrices $B$, $C$, and $\Delta$ to depend on the input sequence. Additionally, its hardware-aware implementation makes Mamba a computationally efficient model, resulting in significant performance gains and resource efficiency. Gu and Dao [24] first proposed the Mamba structure to address the computational burden of long sequences in transformer architecture. They deployed selective State Space Models (SSMs) and a hardware-aware parallel algorithm without using attention or MLP blocks.
Due to its outstanding performance, many researchers have started replacing the Transformer architecture with the Mamba structure. Recently, there have been numerous attempts to adapt the Mamba structure to the medical domain for tasks such as segmentation, classification, and detection, where it has shown superior performance compared to conventional methods. For instance, the authors in [25, 26, 27, 28, 29] proposed Mamba-based methods for medical image segmentation. Yue and Li [30] introduced a medical image classification method that combines convolutional neural networks (CNNs) with the Mamba structure. In their approach, local features are extracted using CNNs, while global features are obtained from the Mamba structure, which are then combined for classifying various medical images. In [31], the author presented MambaMorph, a Mamba-based multi-modality deformable registration framework. Additionally, Schiff et al. [32] utilized bi-directional Mamba blocks to model genomic sequences. Inspired by this pioneering work, we also applied the Mamba structure in our research. Specifically, our goal is to find and match the global features of ICE images to coordinates within a spatial coordinate system, which requires leveraging the entire content of the image. In this context, we believe that the Mamba structure offers more significant performance than Transformers, and we have integrated it into our research.
Figure 2: Overview of the proposed method. (a)Procedure Overview: The clinician first positions the ICE catheter at the home view and then selects the desired viewpoint. The proposed network provides the relative state, which is transformed into robotic coordinates using inverse kinematics (IK). Guided by this information, the clinician controls the ICE catheter, and the entire procedure can be performed iteratively. (b) Iteration step: During the procedure, the user can move sequentially toward the guided target state. To track progress from the start point to the target point, the proposed method acts as a state estimator, continuously updating the current state based on the home view state. Each step can be repeated as needed until the target point is reached.
Figure 3: (a) The constructed cardiac anatomical volume mesh (b) The reconstructed volume mesh for missing structures such as RA, RV, LV. We predicted its center of the mesh based on existing small dataset. (c) The final dataset composition in single subject. There exist several views of the images visualizing different structure in different position. (d) Our training set. We acquired the number of target states based on single image in single subject. The target state has a single position, but different orientation by aligning the fan direction to the center of the target volume.
3 Materials and Methods
3.1 Problem Setup and Notation
The ICE catheter is manually introduced into the femoral vein at the groin through an introducer sheath. A basic ICE study typically begins with the catheter positioned in the mid-right atrium (RA), achieving the home view with the ICE catheter in a neutral position, providing imaging of the RA, tricuspid valve, and right ventricle [33]. Clinicians often rely on this home view as a view that can be easily recovered when encountering difficulties in ICE imaging. We assume this home view as the reference coordinate system for our proposed method.
Since most ICE catheter work is conducted in the RA, our guidance system assumes that there are no obstacles in the RA, and therefore, obstacle avoidance is not addressed in this article. Additionally, we assume that ICE imaging is captured during the diastolic phase, when the heart is relaxed and the ventricles are filled with blood, particularly during atrial systole when the QRS complex is prominent [34].
Finally, while we use the tip’s position and orientation information for modeling, it is important to note that our actual guidance system does not require attaching position/orientation sensors to the ICE catheter. This allows our system to function as a purely ICE-image-based guidance system, eliminating the need for additional sensor attachments on the catheter.
We define the ICE imaging state S = ($x,y,z,\delta x,\delta y,\delta z$) as position and orientation in $\mathbb{SE}(3)$. The ICE catheter has four degrees of freedom: two DOFs for steering the catheter tip in two planes (anterior-posterior knob angle $\theta_{1}$ and the right-left knob angle $\theta_{2}$) using two knobs on the catheter handle, bulk rotation $\theta_{3}$, and translation $d_{4}$ along the major axis of the catheter. Thus, the ICE joint state, J = $(\theta_{1},\theta_{2},\theta_{3},d_{4})$. The ICE imaging, I, is defined as the acquired images, which can be either 2D or 3D.
Since the home view is a neutral position and one of the pre-defined clinical views, we have a transition function from S to J using inverse kinematics ($\mathcal{IK}$) [35] based on the home view (i.e., ${\bf S}_{home}$) as a reference coordinate system:
${}{\bf J}_{home}^{i}=\mathcal{IK}({\bf S}_{home}^{i}).$ (3)
where $i$ represents the index of the possible ICE views.
Then, we can define $\Delta{\bf J}_{(i,j)}$ as follows:
${}\Delta{\bf J}_{(i,j)}=\mathcal{IK}({\bf S}_{home}^{i})-\mathcal{IK}({\bf S}_{home}^{j}),$ (4)
where $i$ and $j$ represent the index of the possible ICE views.
Problem Definition: Given the current imaging ${\bf I}_{curr}$ and a goal index $g$ from a library of views, the objective is to estimate the relative ICE joint state $\Delta{\bf J}_{(curr,g)}$ over time:
$~{}S_{home}^{g}=\mathcal{M}({\bf I}_{home},g),$ (5)
$S_{curr}^{home}=\mathcal{M}({\bf I}_{curr},home),$ (6)
$\Delta{\bf J}_{(curr,g)}=\mathcal{IK}({\bf S}_{home}^{g})-\mathcal{IK}(({\bf S}_{curr}^{home})^{-1}),$ (7)
where $\mathcal{M}$ is our proposed model, which will be detailed in the following sections (Sections 3.3 to 4).
3.2 Procedure Overview
Figure 2 (a) presents an overview of the proposed ICE view guidance system. The procedure is broken down into several key steps, each of which is detailed below:
  1. 1.
    Home view positioning: The procedure begins by positioning the ICE catheter in the mid-right atrium (RA) to achieve the home view. Then, the ICE catheter is placed in a neutral position, which serves as our reference coordinate system.
  2. 2.
    The desired view selection: The clinician selects the goal view (as index $g$) from a predefined library of clinical views. The selected view index is then fed into our model $\mathcal{M}$ to estimate the current imaging state in spatial space $S_{home}^{g}$ (Equation (5)).
  3. 3.
    Compute relative vectors $\Delta{\bf J}_{(curr,g)}$: This process requires the continuous prediction of the relative vector $\Delta{\bf J}_{(curr,g)}$ to provide feedback to the user. To achieve this, it is essential to know the current catheter imaging state ${\bf S}_{curr}^{home}$ (Equation (6)) to calculate the relative vectors needed for movement. A novel aspect of our approach is that our same model also serves as a state estimator for this purpose. Then, using $\mathcal{IK}$, the relative joint vectors $\Delta{\bf J}_{(curr,g)}$ from the current state to the goal state are computed at Equation (7). Figure 2 (b) illustrates the iterative process.
  4. 4.
    Iteration: Repeat the above steps 3-4 until the goal view is successfully reached. Once the goal view is reached, the clinician can repeat the process starting from step 2 to transition from the current view to any other desired view.
Figure 4: The proposed method architecture. (a) The whole network architecture. The center of pink block is MedMamba structure. We added layers to feature mixing with target class code and image. At the output of the network, we added multi-head structure to estimate position and orientation separately. (b) The details of SS2D structure. The image feature is split into sequence in different order and put into SSM model. Then, the output is merged together.
3.3 Dataset
To achieve our goal, we need to model a transition function $\mathcal{M}$. For this, we utilize a dataset of ICE images that are labeled according to the anatomical structures they visualize. These images are paired with corresponding tip position and orientation, which are obtained from the electromagnetic (EM) sensors, providing precise tracking of the catheter’s spatial configuration. Ultimately, the dataset consists of pairs of images with their corresponding anatomical labels and position/orientation information. Additionally, the corresponding volume mesh is required to obtain the target information. For this purpose, we reconstruct volumetric information of cardiac structures using volume contouring algorithms [36]. It generates intra-cardiac images targeting the LA, LAA, Left Inferior Pulmonary Vein (LIPV), Left Superior Pulmonary Vein (LSPV), Right Inferior Pulmonary Vein (RIPV), Right Superior Pulmonary Vein (RSPV), and ESO, as shown in Figure 3 (a). Each dataset contains a target volume mesh with world coordinate system-based positioning, several images visualizing each cardiac part, and the corresponding position and orientation information. We used a dataset collected from the CARTO system (Biosense Webster Inc., USA) [37], comprising data from 858 subjects, with 793 used for training, 25 for validation, and 40 for testing.
3.4 Missing mesh
Our dataset mainly consisted of the LA, LAA, LIPV, LSPV, RIPV, and RSPV volumes. Thus, it includes only a limited dataset for the RA, RV, and LV volumes, comprising five subjects. However, the RV and LV are crucial for intra-cardiac examinations, with the RV view often serving as the home view—the starting point for most procedures.
To overcome this limitation, we estimated the centers of the RA, RV, and LV volumes based on other structures, such as the center points of the LA and LAA, using a subject dataset that includes RA, RV, and LV volumes. This allowed us to determine the relationships between these centers. By applying this relationship to our whole dataset, we estimated the missing centers of the RV, LV, and RA, as illustrated in Figure 3 (b).
3.5 Data pre-processing
In the dataset, each subject has its own coordinate system, known as the world coordinate system, which can vary due to the EM sensors used during data acquisition, leading to inconsistencies across subjects. To address this, we normalized all datasets based on the transducers used in each image, eliminating dependence on the world coordinate system. This normalization is also more practical since transducer manipulation should be performed relative to the transducer itself.
The transducer state in the world coordinate system can be represented as ${\bf S}_{w}^{tr}$ and its corresponding transformation matrix is $T_{w}^{tr}$. We can obtain the target position and orientation relative to the transducer following below equation:
${\bf T}_{tr}^{tar}=({\bf T}_{w}^{tr})^{-1}\cdot{\bf T}_{w}^{tar},$ (8)
where the ${\bf T}_{w}^{tar}$ (${\bf S}_{w}^{tar}$) is target position and orientation, which visualizing each center of the mesh. According to clinical guidance in ICE imaging procedures, the structures are visualized by rotating the ICE, with minimal variation in position. Therefore, we selected a single point inside the RA volume for each target view as target position. To determine the orientation, we aligned the fan direction (target position to center point of fan long axis) to point toward the center of the mesh. The final positions and orientations, ${\bf S}_{tr}^{tar}$, was utilized our target dataset. The composition of the entire dataset for a single subject is shown in Figure 3 (c), while the individual targets for a single subject and single view are shown in Figure 3 (d).
4 Implementation Details
4.1 Network details
We utilized MedMamba as the base model for our method [30]. MedMamba combines convolutional neural networks (CNN) with Mamba, which uses State Space Models (SSM). This combination leverages both local and global information, showing superior performance in medical classification tasks compared to other conventional methods. Consequently, we applied MedMamba to our view navigation research, addressing the regression problem.
The entire architecture is shown in Figure 4. The ICE image, representing the current view, is fed into the network and processed through ResNet50 to extract features [38]. Simultaneously, the target view class, represented as a one-hot vector of size 6 (corresponding to the number of target views), is passed through a single linear layer that adjusts its length to match the squared feature size of the ResNet50 output. This is then resized and concatenated with the image features before being passed through another feature mixing layer composed of a single convolution layer.
MedMamba consists of four sequences of SS-Conv-SSM blocks and patch merging steps, except for the last block, which contains only an SS-Conv-SSM block. The SS-Conv-SSM block splits its feature map into two branches: one for local feature extraction using convolution and the other for global feature extraction using the core element, 2D-Selective-Scan (SS2D). It employs the Cross-Scan Module (CSM), which uses a four-way scanning strategy to achieve a global receptive field without increasing the computational burden. The selective scan state-space model (S6) captures long sequence dependencies from all directional features and merges them to construct the 2D feature map. The patch merging step reduces the spatial dimension to achieve hierarchical representations. At the output of MedMamba, we added two single linear layer heads: one for position regression and the other for orientation regression.
The entire network was trained using quantile loss to estimate the boundaries of estimation. This approach enables the network to estimate not only the median value but also the $q^{th}$ percentile, providing the boundary of estimation with $q^{th}$ reliability. The simplified loss function is shown below:
$l_{quant}(y,\hat{y})=\left\{\begin{array}[]{ll}\alpha\cdot(y-\hat{y}),&\hat{y}\leq y\\ (1-\alpha)\cdot(\hat{y}-y),&\hat{y}>y,\end{array}\right.$ (9)
where the $\alpha$ denote the percentage of $q^{th}$ quantile, y and $\hat{y}$ represent elements of label vector $\mathbf{y}$ and prediction vector $\mathbf{\hat{y}}$, respectively. The total loss function used to train the network is shown below:
$l_{total}=l_{quant}(p,\hat{p})+\lambda*l_{quant}(o,\hat{o}),$ (10)
where the $p,o$ represent the labels for position and orientation, respectively and $\hat{p},\hat{o}$ denote the predicted position and orientation, respectively. The $\lambda$ is weighting parameter.
Figure 5: Validation method. (a) Accuracy calculation: The prediction is considered correct if the predicted view matches the target view, and incorrect otherwise. (b) Distance calculation: The distance is measured from the center of the volume to the predicted view. The normalized distance is calculated by dividing distance with a half distance.
Figure 6: The qualitative results. (a) Randomly selected 12 samples which from specific view visualizing specific structure. The light gray is current view and black is 50 quantile prediction view. Each of dark gray denote the 98, 2 quantile prediction view that represents the boundaries of prediction. (b) Randomly selected 4 samples which from undefined view visualizing non-specific structure.
4.2 Evaluation metric
To evaluate the proposed method, we aimed to determine how accurately the relative position and orientation can guide the fan to the target mesh. We assessed this in two ways. First, we used classification to measure how well the fan’s position and orientation, as predicted by the network, align with the target volume. As shown in Figure 5 (a), a correct view is defined as the fan being properly positioned within the target volume, while a view where the fan is outside the target volume is considered incorrect. We calculated the accuracy of this classification.
Second, we measured the distance between the proposed fan position and the center of the target volume. Specifically, we computed two types of distance factors. As shown in Figure 5 (b), the first type is the real distance, which measures the vertical distance between the center of the target volume and the proposed fan position. The second type is the normalized distance, which standardizes the real distance based on half the size of the volume, where a distance of 1 represents the edge of the volume from its center.
4.3 Training details
The network was trained using the Adam optimizer with a learning rate of 1e-4 during 580 epochs. We employed a batch size of 40, and the entire image dataset was resized to $224\times 224$. In Figure 4, the $N$ is set to [2,2,4,2] for each block, which is default setting of MedMamba-Tiny. The $\lambda$ in Equation 10 is 10 and the $\alpha$ is 0.98 which denotes the $98^{th}$ quantile.The entire framework was implemented in PyTorch and trained on a single A100 GPU.
Figure 7: Quantitative results. The x-axis is distance axis and y-axis is percentage of samples. (a) Total sample distance distribution. (b) Correct sample distance distribution. (c) Correct sample real distance distribution for each target view.
5 Experimental Results
Since the proposed method is designed for real clinical settings and trained on real clinical data, our ability to perform actual validation is limited. Therefore, we present simulation-based results derived from real clinical data to demonstrate the method’s effectiveness. To validate the method, we conducted evaluations using both qualitative and quantitative approaches. The current view image, along with each of the six target view classes, is input into the proposed network, which adjusts the fan’s position and direction based on the prediction values. In the qualitative evaluation, we visually assessed whether the predicted fan was well-contained within the target volume. For the quantitative evaluation, as detailed in Section 4.2, we calculated the success rate percentage of how well the predicted fan was included within the target volume. Additionally, we measured the distance from the center of the target volume to the predicted fan, where a smaller distance indicates a more accurate estimation.
We test our method using 40 subjects which does not overlapped with training subjects. In particular, each subject has several frames, therefore, we test all frames whether it can give desired position and orientation well depending on target view. Finally, 6532 test cases were utilized as final testset.
5.1 Qualitative result
The qualitative results are shown Figure 6. From a current view, we determined the relative position and orientation focusing on a specific view among six different target views. The light gray color indicates the current view position and orientation, while black and dark gray represent the 50 and 98 or 2 percentiles of the predicted fan, respectively.
As shown in the Figure 6 (a), the results include 12 randomly selected cases where the input image corresponds to specific views such as RV, LV, LPV, RPV, LAA, or ESO. In all cases, the predicted fan is clearly contained within the target volume. Particularly, for cases requiring significant movement, such as from LV to RPV, the predictions accurately reflect the desired state. Conversely, for cases requiring minimal movement, such as from LPV to LPV, the predictions also effectively facilitate view navigation. In the Figure 6 (b), it shows 4 randomly selected cases where the input image visualizes undefined views that only intersect the LA volume. In these cases, the proposed method provides accurate target state information for both small and large movements.
5.2 Quantitative result
[…]
The Figure 7 (a) shows the normalized distance between the center of the target volume and the predicted fan. The histogram illustrates the distribution of these distances, including out-of-boundary (OOB) samples. The normalized distance is calculated by dividing the distance from the predicted fan to the center of the target volume by the distance from the center plane to the edge plane. Thus, a distance of 1 represents the edge of the volume, distances below 1 indicate correct predictions, and distances above 1 indicate incorrect predictions. In Figure 7 (b), it shows the histogram for correct sample distances only, with smaller distances corresponding to better predictions. In Figure 7 (c), it gives the correct sample distance distribution for each target view, visualizing histograms for various input images corresponding to target views such as RV, LV, LPV, RPV, LAA, or ESO. For the RV and LV target views, the histograms are slightly noisier due to the smaller number of training samples compared to other target views.
Figure 8: Nearby sample result. ’View A’ and ’View B’ show similar or nearby view of the same area. (a) Both images show LV view (b) Both images show ESO view (c) Both images show LPV view. (d) ’View A’ shows undefined view and ’View B’ shows ESO view. (e) Both images show RPV view.
Table 1: Nearby view prediction difference results
RV LV LPV RPV LAA ESO
Sample 1 5.72[mm] 6.18 [mm] 7.73 [mm] 10.65 [mm] 8.03[mm] 9.95[mm]
(LV view) (0.045,0.121,0.088)[rad] (0.079,0.138,0.105)[rad] (0.079,0.085,0.054)[rad] (0.073,0.064,0.154)[rad] (0.028,0.077,0.090)[rad] (0.075,0.089,0.145)[rad]
Sample 2 3.83[mm] 3.24 [mm] 2.80 [mm] 3.55 [mm] 2.86[mm] 3.33[mm]
(ESO view) (0.038,0.176,0.021)[rad] (0.089,0.065,0.025)[rad] (0.001,0.027,0.072)[rad] (0.082,0.045,0.095)[rad] (0.029,0.013,0.150)[rad] (0.061,0.028,0.051)[rad]
Sample 3 4.12. [mm] 3.92 [mm] 3.59 [mm] 2.96 [mm] 2.92 [mm] 3.19 [mm]
(LPV view) (0.289,0.088,0.234)[rad] (0.095,0.052,0.059)[rad] (0.025,0.079,0.124)[rad] (0.114,0.192,0.050)[rad] (0.073,0.187,0.149)[rad] (0.013,0.067,0.054)[rad]
Sample 4 3.91 [mm] 1.94 [mm] 2.83 [mm] 3.39 [mm] 3.53 [mm] 2.05 [mm]
(Undefined & ESO view) (0.169,0.103,0.466)[rad] (0.089,0.108,0.095)[rad] (0.056,0.002,0.084)[rad] (0.000,0.041,0.098)[rad] (0.040,0.039,0.086)[rad] (0.097,0.017,0.060)[rad]
Sample 5 3.34 [mm] 3.07 [mm] 3.34 [mm] 2.67 [mm] 3.94 [mm] 3.37 [mm]
(RPV view) (0.129,0.010,0.052)[rad] (0.101,0.008,0.017)[rad] (0.173,0.039,0.002)[rad] (0.088,0.022,0.092)[rad] (0.155,0.003,0.011)[rad] (0.144,0.008,0.005)[rad]
5.3 Nearby view result
Since our proposed model is trained based on real clinical data, we evaluated it through a simulation-based process. To assess the robustness of the method, we tested it with views that are similar or nearby to the target viewpoint. Specifically, the method should produce similar results or only minor differences when the current view covers a similar or adjacent area to the target viewpoint. We tested the method using 5 randomly selected samples from different subject that visualized similar or nearby areas.
In the Figure 8, it shows examples of similar or nearby views. In each sample, ’View A’ and ’View B’ visualize very similar areas but from different image. We obtained target states for 6 target views from our proposed method using both images and calculated the differences. The results are summarized in Table 1.
In the case of Sample 1, both views visualize the LV. When targeting the RV viewpoint, the difference between the outputs of the two views is $5.72[mm]$ and $(0.045,0.121,0.088)[rad]$ that can be demonstrated that very similar result. In most of cases, the results show a good result in position and orientation difference for each target view. We demonstrate that the proposed method can account for difference in image to the difference in coordinate.
5.4 State estimation result
As shown in Figure 2 (b), the proposed method can function as a current state estimator. For it to be an effective state estimator, it should provide a stable target state throughout the entire movement from the starting point to the target point. To validate this, we tested our method using a series of images representing movement from the starting point to the target point, visualizing the predicted fan direction for each case throughout the movement. We used four randomly selected subjects and chose a series of images ranging from a arbitrary starting state to a specific target volume.
The results are shown in Figure 9. The colored triangles indicate the fan of the current image, while the black triangles represent the predicted fan. We determined the fan direction by measuring from the predicted position to the center point of the fan’s bottom long axis. This fan direction is separately visualized on the right side of each volume and fan image. Each standard deviation of the predicted position and the center point of the fan’s long axis is calculated and displayed next to the region, highlighted by a light gray elliptical shape.
In Figure 9 (a), we assumed movement from the blue fan point to the RV target. From all intermediate images, we acquired the predicted fan direction, visualized as a black fan. The complete fan direction is displayed on the right side, showing that all predicted fan directions from intermediate images are consistently aligned. In Figure 9 (c), we assumed movement from the pink fan point to the RPV target. The predicted fan directions from all intermediate images are shown on the right side, demonstrating that the proposed method provides a stable target state for all intermediate images during movement and effectively functions as a state estimator.
Figure 9: State estimation results. For several intermediate images during the movement from the start point to the target view, we acquired the target state from the proposed method and determined its fan direction. The left side displays the fan movement and its predicted position, while the right side shows the predicted fan direction vector. (a) RV target movement, (b) LAA target movement, (c) RPV target movement, and (d) LV target movement.
6 Conclusion and Discussion
The intra-cardiac echocardiography (ICE) is an imaging tool used to capture images of the cardiac structures within the cardiac chamber. It is valuable for anatomical reviews, diagnosing congenital heart defects, or assessing valvular heart disease. Additionally, ICE is essential for interventional cardiology, guiding catheter ablation or the closure of septal defects. Despite its importance, effective manipulation of the ICE requires a skilled operator and substantial training due to the difficulty in achieving precise and repeatable controls. To address this issue, We proposed an AI-driven closed-loop view guidance system designed to simplify ICE manipulation based on the current ICE image. Specifically, there exist clinically predefined common views for most procedures, such as RV, LV, LPV, RPV, LAA, and ESO views. Our method provides the relative position and orientation vectors needed to guide the user to their preferred viewpoint from the current view, facilitating easier control of the ICE.
To achieve this, we developed a Mamba-based regression model. The Mamba model, based on the State Space Model (SSM), outperforms traditional transformer architectures. Its selective scan and hardware-aware features enable efficient handling of long-range sequences. Our approach utilizes the MedMamba structure, which combines CNN and Mamba to efficiently capture both global and local features in medical images. For multi-target class regression, we incorporated a feature mixing layer to combine image features with target class codes, and a multi-head structure at the network’s end to separately estimate position and orientation. The network was trained using quantile loss to accurately estimate the boundaries of the predictions. We trained our regression model on a CARTO dataset containing view, position, orientation, and cardiac volume information.
The qualitative and quantitative results demonstrate that our method provides accurate target state information for view visualization. Whether the view is focused on specific structures or not, our method effectively supports view navigation. Additionally, when movement simulation test from the start point to the target state, the intermediate images stably show the target state well, proving that our method functions well as a state estimator.
For future work, we plan to validate our system performance in a beating heart phantom. We believe that our method will simplify ICE manipulation for clinicians and enhance the overall workflow of the procedure.
Disclaimer
The concepts and information presented in this paper are based on research results that are not commercially available. Future availability cannot be guaranteed.